# Multi-height analysis of rational points of toric varieties

### v1

@@ Abstract.
For a quasi-Fano variety $V$ with infinitely many rational points, one may asymptotically study the finite set of rational points with a bounded height $H$ associated with the anticanonical line bundle $\omega_{V}^{-1}$. In [11], [3] and [13], Batyrev, Franke, Manin, Tschinkel and Peyre gave strong evidence supporting conjectures relating the asymptotic behavior of the number of points of bounded height on open subsets of $V$ to geometrical invariants of $V$.
@@
In the case of toric varieties, the asymptotic behavior was studied in [1] and [2] using harmonic analysis on adelic groups. In [18], Salberger was the first to explicitly use universal torsors in relation to points of bounded height. In particular, he was able to give a new proof of the theorem of Batyrev and Tschinkel for smooth projective split toric varieties over $\mathbf{Q}$. His work highlighted how the constant in the conjectural asymptotic formula is related to the product of local densities over universal torsors. This was later studied in a more general setting by Peyre in [14] and [15].
@@
More recently in [17], Peyre proposed a new framework to study the asymptotic distribution of rational points of bounded height. Instead of considering a height relative to one line bundle, it is natural to consider all possible heights. The purpose of this paper is to study the multi-height distribution of rational points of smooth, projective and split toric varieties over $\mathbf{Q}$. We prove that the expected asymptotic behavior holds for these varieties. In particular, we show that there is no accumulating subset for the multi-height distribution in the toric case. This is because we consider the rational points whose multi-height are in the interior of dual of the effective cone.
A result concerning this all heights approach was given in [10], using height zeta functions. It seems that the condition $\beta=\pi^{\vee}(u)$ where $u\in(C_{\operatorname{eff}}(X)^{\vee})^{\circ}$ is missing which would make it possible to construct counterexamples to their main result. Our method using descent is totally different. It is inspired by the work of Salberger in [18] and uses the method of counting the number of points of a lattice in a bounded domain developed by [8].
@@
The author thanks his PhD advisor Emmanuel Peyre for all the remarks and suggestions he made during the writing process of this article.
1. The Manin-Peyre conjecture for the multi-height distribution of rational points on Fano varieties
1.1. Preliminaries
@@
In this paragraph, we recall, following Peyre [17], how to define a system of Arakelov heights for a smooth, projective and geometrically integral variety $V$ over a number field $K$.
@@
If $v\in M_{K}^{0}$, we write $q_{v}=\sharp\mathbf{F}_{v}$ the cardinal of the residual field associated to $v$, while if $v\in M_{K}^{\infty}$, we set $q_{v}=\exp(-[K_{v}:\mathbf{R}])$.
@@
To an inversible sheaf $\mathcal{L}$ over a scheme $T$, we associate the line bundle $L$ whose sections correspond to $\mathcal{L}^{\vee}$. We denote $L^{\times}=L\setminus z(T)$, where $z(T)$ is the closed subscheme given by the zero section.
@@
    If $v\in M_{K}^{0}$, for any $P\in V(K_{v})$, the restriction of $||-||_{v}$ to $E_{P}$ defines an ultrametric norm with values in the image of the absolute value $|-|_{v}$ i.e $q_{v}^{\mathbf{Z}}$.
@@
    If $v$ is a real place, for any $P\in V(K_{v})$, the restriction to $E_{P}$ is an euclidean norm.
@@
    If $v$ is a complex place, for any $P\in V(K_{v})$, the restriction to $E_{P}$ is the square of an hermitian norm.
@@
    There exists a finite number of places $S\subset M_{K}$ (which contains $M_{K}^{\infty}$) and a model $\mathcal{E}\rightarrow\mathcal{V}$ of $E\rightarrow V$ over $\mathcal{O}_{S}$ such that for any place $v\in M_{K}\setminus S$, for any $P\in\mathcal{V}(\mathcal{O}_{v})$,
    $\mathcal{E}_{P}=\{y\in E_{P}|\ \ ||y||_{v}\leqslant 1\}$
@@
For smooth, projective and split toric varieties over $\mathbf{Q}$ such model always exists over $\mathbf{Z}$.
The following example is particularly important, because, as we shall later see, the natural system of heights over toric varieties can be deduced from this example.
@@
This defines an adelic norm $(||-||_{v})_{v\in M_{K}}$ over the line bundle $\mathcal{O}_{\mathbf{P}^{N}}(-1)\rightarrow\mathbf{P}^{N}$ which corresponds to the $\mathbf{G}_{m}$-torseur $\mathbf{A}^{N+1}\setminus\{0\}\rightarrow\mathbf{P}^{N}$.
@@
If $L\rightarrow V$ is a line bundle, $\{||-||_{v}\}$ an adelic norm over $L$, then we define the height $H_{L}:V(K)\rightarrow\mathbf{R}_{+}$ for any $x\in V(K)$ such that :
$H_{L}(x)=\prod\limits_{v\in M_{K}}||y||_{v}^{-1}$
for $y\in L^{\times}(K)$ over $x$.
@@
Let $L$ and $L^{\prime}$ be two adelic line bundles over $V$ (i.e equipped with an adelic norm). Let $(||-||_{v})$ be the adelic norm over $L$.
@@
We denote by $\mathcal{H}(V)$ the set of equivalence classes of adelic line bundles. This set is naturally equipped with a structure of group given by the tensorial product of adelic line bundle.
Hence, we can define the notion of system of heights and of multi-heights over $V$.
@@
$\mathrm{o}:\mathcal{H}(V)\rightarrow\operatorname{Pic}(V)$
@@
such that for any $P\in V(K)$, $h(P)$ is the linear form defined by $L\in\operatorname{Pic}(V)$,
$h(P)(L)=\log H_{L}(P)$
In order to state the conjecture made by Peyre [17, question 4.8], we shall define a measure $\nu$ on $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ as follows:
@@
[…]
for any a compact subset $\mathrm{D}$ of $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$, where the Haar measure $dy$ on $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ is normalised so that the covolume of the dual lattice of the Picard group is one.
@@
We shall say that $V$ is quasi-Fano if it verifies the following hypothesis:
@@
In particular, any Fano variety and any smooth, projective and split toric varieties are quasi-Fano. With these notations, Peyre states the following conjecture:
@@
$\mathrm{D}_{B}=\mathrm{D}_{1}+\log(B)u$
There exists a ’small’ subset $T\subset V(K)$ so that we have an asymptotic behaviour of the form:
@@
where $\beta(V)=\sharp H^{1}(\mathbf{Q},\operatorname{Pic}(V_{\overline{\mathbf{Q}}}))$ and $\tau(V)$ is the Tamagawa number of $V$.
The Tamagawa number of $V$, $\tau(V)$ is defined in 2.20.
@@ 1.2. The Manin-Peyre conjecture in the case of toric varieties
In this article all toric varieties are supposed to be smooth, projective and split toric variety over $\mathbf{Q}$. Let $X$ be such toric variety. We shall use in the rest of this paper that the anticanonical line bundle $\omega_{X}^{-1}$ is big, i.e is in the interior of the effective cone $\text{C}_{\text{eff}}(X)$.
@@
We denote by $T\subset X$ its open torus, by $\Sigma$ a fan associated to $X=X_{\Sigma}$, $\Sigma(1)$ the set of edges of $\Sigma$ and by $m$ the number of edges. We denote by $t$ the rank of the Picard group $\operatorname{Pic}(X)$. With our hypothesis, it is a free abelian group of finite rank.
To an edge $\rho\in\Sigma(1)$, we associate the $T$-invariant divisor $D_{\rho}$. Recall that
$\omega_{X}^{-1}=\sum\limits_{\rho\in\Sigma(1)}[D_{\rho}]$
@@
the natural map which associates to a $T$-invariant divisor $D$ the class of the associated line bundle in $\operatorname{Pic}(X)$ which we write $[D]=\pi(D)$.
We denote by $N=X_{*}(T)$ the group of subgroups of one parameters of $T$ and by $M=X^{*}(T)$ its dual, the group of characters of $T$.
If $L$ is a line bundle over $X$, there exists a natural way to define an adelic norm $(||-||_{v})_{v\in M_{K}}$ on $L$, following [18, definition 9.2] (see also [3, proposition 2.1.2]). Let us explain how one can interpret this adelic norm in the case of $\mathbf{P}^{N}$.
@@
For $0\leq i\leq N$, we denote by $D_{i}$ the divisor $\mathbf{G}_{m}^{N}$-invariant of $\mathbf{P}^{N}$ corresponding to the hyperplane $\{X_{i}=0\}\subset\mathbf{P}^{N}$.
Then the adelic norm over $\mathcal{O}_{\mathbf{P}^{N}}(1)$ associated to $D_{i}$ corresponds to the one of the previous example 1.5.
Proof.
This is a consequence of [18, proposition 9.8]. ∎
Remark 1.12.
We can moreover observe that the construction of the adelic norm in [18, definition 9.2] over $\mathcal{O}(D)$ depends only on the class of the $T$-invariant divisor $D$ modulo $M=X^{*}(T)$, that is to say its class $[D]$ in $\operatorname{Pic}(X)$.
Proposition 1.13.
The heights system $s_{X}:\operatorname{Pic}(X)\rightarrow\mathcal{H}(X)$ defined on a toric variety $X$ is functorial for the morphisms of toric varieties.
[…]
We shall later see using versal torsor that this way of defining a height system over our toric variety $X$ (and a representative of its heights system) makes the count of rational points of our variety easier.
@@
Let $f$ be the smallest positive integer such that there exists $f$ rays of $\Sigma$ not contained in a cone of $\Sigma$. Note that $f\geqslant 2$.
@@
Let $\varepsilon\in]0,1-\frac{1}{f}[$. Then we have an asymptotic behaviour of the form:
$\sharp(X(\mathbf{Q}))_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{=}\nu(\mathrm{D}_{1}).\tau(X).B^{\langle\omega_{X}^{-1},u\rangle}\left(1+O\left(B^{-(1-\frac{1}{f}-\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}\right)\right)$
where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 1.8.
@@
If $\mathbf{b}=(L_{1},..,L_{r})$ is a family of line bundles such that there classes $\mathbf{R}$-generate the Picard group $\operatorname{Pic}(X)_{\mathbf{R}}$, $\alpha_{1},..,\alpha_{r},\beta_{1},..,\beta_{r}$ are real numbers such that for any $i$, $\alpha_{i}<\beta_{i}$, then we define a compact polyhedron of $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$:
@@ 2. The notion of universal torsors
In this section, we shall present how the study of the multi-height distribution of rational points of toric varieties naturally lifts to a counting problem in the universal torsor at the origin.
2.1. Generalities about versal torsors

### v2

@@ 1. Introduction
For a quasi-Fano variety $V$ with infinitely many rational points, one may asymptotically study the finite set of rational points with a bounded height $H$ associated with the anticanonical line bundle $\omega_{V}^{-1}$. In [12], [1] and [15], Batyrev, Franke, Manin, Tschinkel and Peyre gave strong evidence supporting conjectures relating the asymptotic behavior of the number of points of bounded height on open subsets of $V$ to geometric invariants of $V$.
@@
In the case of toric varieties, the asymptotic behavior was studied in [2] and [3] using harmonic analysis on adelic groups. In [20], Salberger was the first to explicitly use universal torsors to study points of bounded height. In particular, he was able to give a new proof of the theorem of Batyrev and Tschinkel for smooth projective split toric varieties over $\mathbf{Q}$. His work highlighted how the constant in the conjectural asymptotic formula is related to the product of local densities on universal torsors. This was later studied in a more general setting by Peyre in [16] and [17].
@@
More recently in [19], Peyre proposed a new framework to study the asymptotic distribution of rational points with bounded height. Instead of considering a height relative to one line bundle, it is natural to consider all possible heights. The purpose of this paper is to study the multi-height distribution on rational points of smooth, proper and split toric varieties over $\mathbf{Q}$. We prove that the expected asymptotic behavior holds for these varieties. In particular, we show that there is no accumulating subset for the multi-height distribution in the toric case. This is because we consider the rational points whose multi-height is in the interior of the dual of the effective cone.
A result concerning this all heights approach was given in [11], using height zeta functions. However, the condition $\beta=\pi^{\vee}(u)$ where $u\in(C_{\operatorname{eff}}(X)^{\vee})^{\circ}$ seems to be missing in [11, theorem 1.1]. If one considers the case of the projective line $\mathbf{P}^{1}$ with $(\beta_{1},\beta_{2})=(1,2)$, their result would imply that the asymptotic behavior is of the form $C.B^{3}$ whereas it is known to be of the form $C.B^{2}$.
In any case, our method using descent is totally different. It is inspired by the work of Salberger in [20] and uses the method of counting the number of points of a lattice in a bounded domain developed in [9] by Davenport.
Structure. We begin in Section 2 by recalling the multi-height formalism and the conjecture on the asymptotic distribution of rational points for quasi-Fano varieties associated with this formalism. We then state the main result established in this article for toric varieties. In the following section, we introduce a crucial concept for proving our result: the universal torsor. We explain how to lift the height of a point locally to the universal torsor, describe the specific case of toric varieties, and conclude by linking the measure on universal torsors with the Tamagawa number of the variety. In the final section, we establish our result using estimation techniques from the geometry of numbers and the Möbius inversion formula.
@@
Acknowledgment. I would like to thank my PhD advisor Emmanuel Peyre for all the remarks and suggestions he made during the writing process of this article. I would also like to thank the referee for their help in improving the exposition.
2. Terminology and the precise statement
2.1. Arakelov heights
@@
In this paragraph, we recall, following Peyre [19], how to define a system of Arakelov heights for a smooth, proper and geometrically integral variety $V$ over a number field $K$.
@@
If $v\in M_{K}^{0}$, we write $q_{v}=\sharp\mathbf{F}_{v}$ the cardinal of the residual field associated to $v$, while if $v\in M_{K}^{\infty}$, we set $q_{v}=\exp([K_{v}:\mathbf{R}])$.
@@
To an invertible sheaf $\mathcal{L}$ over a scheme $T$, we associate the line bundle $L$ whose sections correspond to $\mathcal{L}$. We denote $L^{\times}=L\setminus z(T)$, where $z(T)$ is the closed subscheme given by the zero section.
@@
    If $v\in M_{K}^{0}$, for any $P\in V(K_{v})$, the restriction of $||-||_{v}$ to $E_{P}$ defines an ultrametric norm with values in the image of the absolute value $|-|_{v}$ i.e. $q_{v}^{\mathbf{Z}}$.
@@
    If $v$ is a real place, for any $P\in V(K_{v})$, the restriction to $E_{P}$ is a Euclidean norm.
@@
    If $v$ is a complex place, for any $P\in V(K_{v})$, the restriction to $E_{P}$ is the square of a Hermitian norm.
@@
    There exists a finite set of places $S\subset M_{K}$ (which contains $M_{K}^{\infty}$) and a model $\mathcal{E}\rightarrow\mathcal{V}$ of $E\rightarrow V$ over $\mathcal{O}_{S}$ such that for any place $v\in M_{K}\setminus S$, for any $P\in\mathcal{V}(\mathcal{O}_{v})$,
    $\mathcal{E}_{P}=\{y\in E_{P}|\ \ ||y||_{v}\leqslant 1\}.$
@@
For smooth, proper and split toric varieties over $\mathbf{Q}$ such models always exists over $\mathbf{Z}$.
The following example is particularly important, because, as we shall later see, a natural system of heights over toric varieties can be deduced from this example.
@@
This defines an adelic norm $(||-||_{v})_{v\in M_{K}}$ on the line bundle $\mathcal{O}_{\mathbf{P}^{N}}(-1)\rightarrow\mathbf{P}^{N}$ which corresponds to the $\mathbf{G}_{m}$-torseur $\mathbf{A}^{N+1}\setminus\{0\}\rightarrow\mathbf{P}^{N}$.
@@
If $L\rightarrow V$ is a line bundle, $\{||-||_{v}\}$ an adelic norm on $L$, then we define the height $H_{L}:V(K)\rightarrow\mathbf{R}_{+}$ for any $x\in V(K)$ by:
[…]
where $y\in L^{\times}(K)$ lies above $x$.
@@
Let $L$ and $L^{\prime}$ be two adelic line bundles over $V$ (i.e. equipped with an adelic norm). Let $(||-||_{v})$ be the adelic norm on $L$.
@@
We denote by $\mathcal{H}(V)$ the set of equivalence classes of adelic line bundles. This set is naturally equipped with a structure of group given by the tensorial product of adelic line bundles.
Hence, we can define the notion of system of heights and of multi-height over $V$.
@@
$\mathrm{o}:\mathcal{H}(V)\rightarrow\operatorname{Pic}(V).$
@@
such that for any $P\in V(K)$, $h(P)$ is the linear form defined for any $L\in\operatorname{Pic}(V)$ by:
$h(P)(L)=\log H_{s(L)}(P).$
Remark 2.8.
The multi-height is well-defined as $h(P)(L)$ does not depend on the choice of the representative of the equivalence class (see [19, remark 4.2]).
2.2. The multi-height conjecture
In order to state the conjecture made by Peyre [19, question 4.8], we shall define a measure $\nu$ on $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ as follows:
@@ 2.2. The multi-height conjecture
$\nu(\mathrm{D}):=\int\limits_{\mathrm{D}}e^{\langle\omega_{V}^{-1},y\rangle}dy$
for any compact subset $\mathrm{D}$ of $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$, where the Haar measure $dy$ on $\operatorname{Pic}(V)^{\vee}_{\mathbf{R}}$ is normalized so that the covolume of the dual lattice of the Picard group is one.
@@
We shall say that $V$ is quasi-Fano if it satisfies the following hypotheses:
@@
Any Fano variety and any smooth, proper and split toric variety is quasi-Fano. Let us quickly explain where the last claim comes from. The effective cone of such a variety is generated by the divisors associated to each edge of the fan defining the toric variety, and its anticanonical bundle is the sum of these divisors, so it is in the interior of the effective cone. Hence we get that the first and the fifth hypotheses are verified. As such a variety is covered by open subschemes isomorphic to $\mathbf{A}^{d}$, where $d$ is the dimension of the toric variety, we get the second assertion and the fact that its geometrical Brauer group is trivial. For such toric variety, by [8, proposition 4.2.5], the geometric Picard group is torsion free. Finally, by [8, theorem 9.2.3], the third hypothesis is verified. Peyre conjectures the following in [19, question 4.8]:
@@
$\mathrm{D}_{B}=\mathrm{D}_{1}+\log(B)u.$
There exists an exceptional subset $T\subset V(K)$ such that we have an asymptotic behaviour of the form:
@@
where $\beta(V)=\sharp H^{1}(\mathbf{Q},\operatorname{Pic}(V_{\overline{\mathbf{Q}}}))$ and $\tau(V)$ is the Tamagawa number of $V$ (see definition 3.22).
@@ 2.3. The Manin-Peyre conjecture in the case of toric varieties
Let $X$ be a smooth, proper and split toric variety over $\mathbf{Q}$. We shall use in the rest of this paper that the anticanonical line bundle $\omega_{X}^{-1}$ is big, i.e. is in the interior of the effective cone $\text{C}_{\text{eff}}(X)$.
@@
We denote by $T\subset X$ its open torus, by $\Sigma$ a fan associated to $X=X_{\Sigma}$, by $\Sigma(1)$ the set of edges of $\Sigma$ and by $m$ the number of edges. We denote by $t$ the rank of the Picard group $\operatorname{Pic}(X)$. With our hypothesis, it is a free abelian group of finite rank. To an edge $\rho\in\Sigma(1)$, we associate the $T$-invariant divisor $D_{\rho}$.
@@
the natural map which associates to a $T$-invariant divisor $D$ the class of the associated line bundle in $\operatorname{Pic}(X)$ which we write $D]=\pi(D)$. Recall that
$\omega_{X}^{-1}=\sum\limits_{\rho\in\Sigma(1)}[D_{\rho}].$
We denote by $N=X_{*}(T)$ the 1-parameter subgroups of $T$ and by $M=X^{*}(T)$ its dual, the group of characters of $T$.
Convention 2.12.
A line bundle $L$ over $X$ is endowed with the associated natural adelic norm $(||-||_{v})_{v\in M_{K}}$ following [20, definition 9.2] (see also [1, proposition 2.1.2]).
Let us explain how one can interpret this adelic norm in the case of $\mathbf{P}^{N}$.
Example 2.13.
For $0\leq i\leq N$, we denote by $D_{i}$ the divisor $\mathbf{G}_{m}^{N}$-invariant of $\mathbf{P}^{N}$ corresponding to the hyperplane $\{X_{i}=0\}\subset\mathbf{P}^{N}$. Then the adelic norm over $\mathcal{O}_{\mathbf{P}^{N}}(1)$ associated to $D_{i}$ corresponds to the one of the example 2.5.
Proof.
This is a consequence of [20, proposition 9.8]. ∎
@@
We can moreover observe that the construction of the adelic norm in [20, definition 9.2] over $\mathcal{O}(D)$ depends only on the class of the $T$-invariant divisor $D$ modulo $M=X^{*}(T)$, that is to say its class $[D]$ in $\operatorname{Pic}(X)$.
We shall later see using universal torsor that this way of defining a height system over our toric variety $X$ (and a representative of its system of heights) makes the count of rational points of our variety easier.
@@
Let $l$ be the smallest positive integer such that there exists $l$ rays of $\Sigma$ not contained in a cone of $\Sigma$. Note that $l\geqslant 2$.
@@
Let $\varepsilon\in]0,1-\frac{1}{l}[$. Then we have an asymptotic behaviour of the form:
$\sharp(X(\mathbf{Q}))_{h\in\mathrm{D}_{B}}\underset{B\rightarrow+\infty}{=}\nu(\mathrm{D}_{1}).\tau(X).B^{\langle\omega_{X}^{-1},u\rangle}\left(1+O\left(B^{-(1-\frac{1}{l}-\varepsilon).\min\limits_{\rho\in\Sigma(1)}\langle[D_{\rho}],u\rangle}\right)\right)$
where $\nu$ is the measure over $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$ defined as in 2.9.
Remark 2.16.
In forthcoming work, we will see that, provided that we restrict ourselves to a generating family of $\operatorname{Pic}(X)$ consisting of effective line bundles, it is possible to obtain a similar result without imposing a lower bound on the height, but this introduces a $\log(B)$ factor in the error term. We hope also to show that this result implies the usual form of Manin’s conjecture.
@@
If $\mathbf{b}=(L_{1},..,L_{r})$ is a family of line bundles such that their classes $\mathbf{R}$-generate the Picard group $\operatorname{Pic}(X)_{\mathbf{R}}$, $\alpha_{1},..,\alpha_{r},\beta_{1},..,\beta_{r}$ are real numbers such that for any $i$, $\alpha_{i}<\beta_{i}$, then we define a compact polyhedron of $\operatorname{Pic}(X)^{\vee}_{\mathbf{R}}$:
@@ 3. The notion of universal torsors
In this section, we shall present how the study of the multi-height distribution of rational points of toric varieties naturally lifts to a counting problem in the universal torsor.
3.1. Generalities about universal torsors